What is the maximum you can finance if you can only afford a monthly payment
of $200?   The interest rate is 5%. The term is three years.
Answer #1

Using financial calculator , to find the value of the finance

Inputs: Fv = 0

N = 3× 12 =36

I/Y = 5% / 12 = 0.4167

Pmt = -200

Pv = compute

We get, Pv = $6,673.14

Therefore the maximum finance which can be afforded is $6,673.14